OpenStudy (anonymous):

1+sinθ-cosθ/1+sinθ+cosθ=tanθ/2

OpenStudy (anonymous):

why are thetas being displayed as ???

OpenStudy (anonymous):

\[1+\sin \alpha-\cos \alpha/1+\sin \alpha+\cos \alpha=\tan \alpha/2\]

OpenStudy (anonymous):

Please give brackets and write the question, is it to prove lhs =rhs ??

OpenStudy (anonymous):

I have got his question.. :)

OpenStudy (anonymous):

\[\frac{1 + \sin(x) - \cos(x)}{1 + \sin(x) + \cos(x)} = \tan(\frac{x}{2})\]

OpenStudy (dumbcow):

it may help to make u = 1+sin and the half angle identity for tangent is given as (1-cos)/sin \[\rightarrow \frac{u-\cos x}{u+ \cos x} = \frac{1- \cos x}{\sin x}\] multiply left side by conjugate expand and simplify until both sides are equal

OpenStudy (anonymous):

Multiply by its conjugate after combining the terms as: \[\frac{(1 + \sin(x) - \cos(x)) \times ( 1 - (\sin(x) + \cos(x))}{(1 + (\sin(x) + \cos(x))) \times (1 - (\sin(x) + \cos(x)))}\]

OpenStudy (anonymous):

Sorry, not conjugate, it is just like rationalizing the denominator..

OpenStudy (anonymous):

The denominator simplifies to \(2 \sin(x) \cos(x)\) or simply \( \sin(2x)\)..

OpenStudy (anonymous):

And in numerator when you will multiply the two big brackets, they will simplify to : \[2 \cos(x) - 2 \cos^2(x)\]

OpenStudy (anonymous):

Thereby giving you: \[\implies \frac{2 \cos(x) - 2 \cos^2(x)}{2 \sin(x) \cos(x)} \implies \frac{1 - \cos(x)}{\sin(x)}\]

ganeshie8 (ganeshie8):

you could also try \[\begin{align}\frac{1 + \sin(x) - \cos(x)}{1 + \sin(x) + \cos(x)} &= \dfrac{2\sin^2(x/2)+2\sin(x/2)\cos(x/2)}{2\cos^2(x/2) + 2\sin(x/2)\cos(x/2)} \\~\\ &= \dfrac{\sin(x/2)[2\sin(x/2) + 2\cos(x/2)]}{\cos(x/2)[2\cos(x/2) + 2\sin(x/2)]}\\~\\ &=\dfrac{\sin(x/2)}{\cos(x/2)} \end{align}\]